Write the slope-intercept equation for the line that passes through (-1, -3) and (-2, 5). Please show all of your work
Answer by mananth(16949)

x1----y1----x2----y2
-1-----3-----2----5

slope m =(y2-y1)/(x2-x1)
m=-8

Plug value of the slope and point (-1,-3) in
Y =mx+b
-3.00=8+ b
b=-3-8
b=-11
So the equation will be
Y =-8x-11